### v1

# Sums of squares of integers from residue classes

Abstract.
A subset $\mathcal{A}\subseteq\mathbb{Z}$ is called $s$-almost square universal if every sufficiently large positive integer can be written as a sum of at most $s$ squares of integers from $\mathcal{A}$. In this article, we study the minimal number $\mathrm{ASU}(\mathcal{A}_{d,m})$ with this property, where $\mathcal{A}_{d,m}$ denotes the residue class of $d$ modulo $m$, with $m\in\mathbb{N}$ and $d\in\mathbb{Z}$. We further prove that $\mathcal{A}_{d,m}$ is $s$-square universal for some $s\in\mathbb{N}$ if and only if $d\equiv\pm 1\pmod{m}$, and determine the minimal such number $\mathrm{SU}(\mathcal{A}_{d,m})$ in these cases.
Key words and phrases:
Sums of squares, Almost square universality, Square Universality
2020 Mathematics Subject Classification:
11E12, 11E20, 11E25
This research was supported by the College of Education, Korea University Grant in 2024.
1. Introduction
Waring’s problem asks whether, for a given integer $k\geq 2$, there exists a number $r$ such that every positive integer is a sum of at most $r$ $k$-th powers of positive integers. In 1770, the same year that Edward Waring posed his conjecture, Lagrange proved the Four Square Theorem, which states that every positive integer is the sum of at most four squares of integers. This theorem, though proved independently, resolved Waring’s problem in the special case $k=2$. In general, the existence of such a number was proved by Hilbert [12] in 1909, and the determination of the smallest such number $g(k)$ has been extensively studied.
Building upon this, a closely related and indeed well-studied refinement considers the smallest number $G(k)$ such that every sufficiently large integer can be represented as a sum of at most $G(k)$ $k$-th powers of positive integers. For $k=2$, Legendre’s Three Square Theorem implies that no positive integer of the form $4^{c}(8d+7)$ for integers $c,d$, can be expressed as a sum of three squares of integers. Thus, $g(2)=G(2)=4$. For more on these invariants $g(k)$ and $G(k)$, we refer the reader to the survey notes [10] and [21], and the references therein.
In what follows, we focus on the case $k=2$, that is, sums of squares. This naturally leads to the study of positive definite integral quadratic forms and their universality properties. A classical example is Ramanujan’s classification [19] of 55 diagonal quaternary quadratic forms that are universal—that is, quadratic forms representing all positive integers. Dickson [8] refined this list by removing the form $x^{2}+2y^{2}+5z^{2}+5w^{2}$, which fails to represent $15$. More recently, Conway and Schneeberger proved the celebrated 15-Theorem, which gives a complete criterion for universality: a positive definite integral quadratic form is universal if and only if it represents the integers $1,2,3,5,6,7,10,14,$ and $15$. Bhargava [2] later gave a simplified proof and, together with Hanke [3], extended the result to the 290-Theorem for integer-valued quadratic forms.
Beyond classical universality results, there has been considerable interest in representations by sums of squares with restricted variables, such as restrictions to almost primes. Brüdern and Fouvry [5] proved that every sufficiently large integer $n\equiv 4\ (\mathrm{mod}\ 24)$ can be expressed as a sum of four squares of integers, each having at most 34 prime factors, and Blomer and Brüdern [4] obtained a corresponding result for three squares.
Motivated by these developments, a natural generalization is to consider sums of squares drawn from specific subsets $\mathcal{A}$$\subseteq{\mathbb{Z}}$, leading to a Waring-type problem. We say $\mathcal{A}$} is $s$-square universal (resp. $s$-almost square universal) if every positive integer (resp. every sufficiently large integer) can be written as a sum of at most $s$ squares from ${\mathcal{A}}^{2}$. The minimal such $s$, if it exists, is denoted by $\mathrm{SU}({\mathcal{A}})$ (resp. $\mathrm{ASU}({\mathcal{A}})$).
For example, in [15], this problem was studied for ${\mathcal{A}}={\mathbb{Z}}\setminus p{\mathbb{Z}}$ with each prime $p$, where it is shown that
$\mathrm{SU}({\mathbb{Z}}\setminus 2{\mathbb{Z}})=10,\ \mathrm{SU}({\mathbb{Z}}\setminus 3{\mathbb{Z}})=6,\ \mathrm{SU}({\mathbb{Z}}\setminus 5{\mathbb{Z}})=5,\text{ and }\mathrm{SU}({\mathbb{Z}}\setminus 2{\mathbb{Z}})=4\text{ for any prime }p\geq 7.$
Although these stated results concern square universality, their proofs in fact establish the corresponding almost square universality results as well:
$\mathrm{ASU}({\mathbb{Z}}\setminus 2{\mathbb{Z}})=10,\ \mathrm{ASU}({\mathbb{Z}}\setminus 3{\mathbb{Z}})=6,\ \mathrm{ASU}({\mathbb{Z}}\setminus p{\mathbb{Z}})=4\text{ for any prime }p\geq 7.$
More recently, the case ${\mathcal{A}}=S_{\rho}:={\mathbb{Z}}\setminus\{\pm\rho\}$ with $\rho\in{\mathbb{N}}$ was studied in [6]. Regarding almost square universality, it is shown that $\mathrm{ASU}(S_{\rho})=4$ for all $\rho\in{\mathbb{N}}$. For square universality, one may observe that $\mathrm{SU}(S_{1})$ does not exist, since $S_{1}$ excludes $\{\pm 1\}$. For other cases, it was proved that
[…]
In this article, we study the above concepts for the set ${\mathcal{A}}={\mathcal{A}}_{d,m}$, the residue class of $d$ modulo $m$, where $m\in{\mathbb{N}}$ and $d\in{\mathbb{Z}}$. Under the natural assumption that $m$ and $d$ are relatively prime—since otherwise $\mathrm{ASU}(\mathcal{A}_{d,m})$ cannot exist—we obtain the following main result characterizing almost square universality.
Theorem 1.1.
For $m\in{\mathbb{N}}$ and $d\in{\mathbb{Z}}$ with $(m,d)=1$, we have
[…]
Moreover, we characterize when ${\mathcal{A}}_{d,m}$ is $s$-square universal for some $s\in{\mathbb{N}}$, and determine the exact value of $\mathrm{SU}(\mathcal{A}_{d,m})$ in those cases.
Theorem 1.2.
The set ${\mathcal{A}}_{d,m}$ is $s$-square universal for some $s\in{\mathbb{N}}$ if and only if $d\equiv\pm 1\ (\mathrm{mod}\ m)$. Furthermore, for such integers $d$, we have
[…]
This article is organized as follows. In Section 2, we review preliminary results for quadratic spaces, lattices, and lattice cosets. Section 3 is devoted to proving almost square universality of ${\mathcal{A}}_{d,m}$ and Theorem 1.1. Finally, Section 4 addresses the proof of square universality, establishing Theorem 1.2.
2. Preliminaries
In this section, we introduce notation and terminology of quadratic spaces, lattices, and lattice cosets, and their theta series.
2.1. Quadratic spaces and lattices
A quadratic space over $\mathbb{Q}$} is a triple $(V,B,Q)$ of a $\mathbb{Q}$}-vector space $V$ with a quadratic map $Q:V{\,\rightarrow\,}{\mathbb{Q}}$ and the associated bilinear form $B:V\times V{\,\rightarrow\,}{\mathbb{Q}}$ defined by $B(x,y)=\frac{1}{2}(Q(x+y)-Q(x)-Q(y))$. We say that $V$ is positive definite if $Q(x)>0$ for all nonzero elements $x\in V$. Throughout this article, we assume that all quadratic spaces $V$ are positive definite.
A quadratic $\mathbb{Z}$}-lattice on $V$ is a free $\mathbb{Z}$}-module $L={\mathbb{Z}}e_{1}+\cdots+{\mathbb{Z}}e_{r}$ of rank $r$ such that ${\mathbb{Q}}L=V$. The $r\times r$ symmetric matrix $M_{L}:=B(e_{i},e_{j})$ is called the Gram matrix of $L$ with respect to a basis $e_{1},\ldots,e_{r}$, and we write $L\cong M_{L}$. If the Gram matrix $M_{L}$ is diagonal, we simply write $L\cong\langle Q(e_{1}),\ldots,Q(e_{r})\rangle$. The quadratic form associated to $L$ is defined by $Q_{L}(x_{1},\ldots,x_{r}):=\bm{x}^{t}M_{L}\bm{x}$ where $\bm{x}=(x_{1},\ldots,x_{r})^{t}$. The discriminant $d_{L}$ of $L$ is the determinant of the Gram matrix $M_{L}$, and the level $N_{L}$ is defined to be the smallest positive integer $N$ such that $NM_{L}^{-1}$ has coefficients in $\mathbb{Z}$}. We define the localization of $V$ and $L$ at any prime spot $p$ by $V_{p}:=V\otimes_{\mathbb{Q}}{\mathbb{Q}}_{p}$ and $L_{p}=L\otimes_{\mathbb{Z}}{\mathbb{Z}}_{p}$.
For two $\mathbb{Z}$}-lattices $\ell$ and $L$, we say that $\ell$ is represented by $L$ if there is a linear map $\sigma:\ell{\,\rightarrow\,}L$ such that $B(\sigma x,\sigma y)=B(x,y)$ for all $x,y\in\ell$. Such a linear map is called a representation from $\ell$ to $L$, and we write $\ell{\,\rightarrow\,}L$ if there exists one. If $\ell{\,\rightarrow\,}L$ and $L{\,\rightarrow\,}\ell$, then we say that they are isometric, and write $\ell\cong L$.
These notations extend naturally to ${\mathbb{Z}}_{p}$-lattices $\ell_{p}$ and $L_{p}$. We say that $\ell$ is locally represented by $L$ if $\ell_{p}{\,\rightarrow\,}L_{p}$ for all primes $p$. The genus $\text{gen}(L)$ of $L$ is defined by
$\text{gen}(L):=\{K\text{ on }{\mathbb{Q}}L\mid K_{p}\cong L_{p}\text{ for all primes }p\}.$
The isometry induces an equivalence relation on $\text{gen}(L)$, and the number of equivalence classes is called the class number of $L$.
It is well-known that if $\ell$ is locally represented by $L$, then there is a $\mathbb{Z}$}-lattice $L^{\prime}\in\text{gen}(L)$ such that $\ell{\,\rightarrow\,}L^{\prime}$ (see [17, 102:5]). Consequently, if $L$ is of class number one, then every locally represented $\ell$ is also globally represented by $L$.
Utilizing this fact together with a detailed study of the representation of binary $\mathbb{Z}$}-lattice $\ell\cong\left(\begin{smallmatrix}4&b\\ b&a\end{smallmatrix}\right)$ by the quaternary $\mathbb{Z}$}-lattice $L\cong\langle 1,1,1,1\rangle$, a generalization of Cauchy’s lemma was established in [14]. This lemma plays a crucial role in our arguments. For earlier versions of related results, see [16].
Lemma 2.1.
For $a,b\in{\mathbb{Z}}$, the system of equations
$\begin{cases}x_{1}^{2}+x_{2}^{2}+x_{3}^{2}+x_{4}^{2}=a\\ x_{1}+x_{2}+x_{3}+x_{4}=b\end{cases}$ (2.1)
has a solution $x_{1},x_{2},x_{3},x_{4}\in{\mathbb{Z}}$ if and only if
$a\equiv b\ (\mathrm{mod}\ 2),\quad 4a-b^{2}\geq 0\quad\text{and}\quad 4a-b^{2}\neq 4^{\alpha}(8\beta+7)\text{ for some $\alpha,\beta\in{\mathbb{Z}}$}.$ (2.2)
Proof.
Since the case when $4a-b^{2}>0$ has been proved in [14, Lemma 4.2], let us assume that $4a-b^{2}=0$. First, $b$ should be even since $4a=b^{2}$. Hence $a$ is also even as $a\equiv b\ (\mathrm{mod}\ 2)$. Writing $a=2a^{\prime}$ and $b=2b^{\prime}$, we have $2a^{\prime}=(b^{\prime})^{2}$, so $b^{\prime}=2b^{\prime\prime}$ for some $b^{\prime\prime}\in{\mathbb{Z}}$ and $a=(b^{\prime\prime})^{2}$. Thus, $x_{i}=(b^{\prime\prime})^{2}$ is a solution to (2.1). ∎
2.2. Quadratic lattice cosets
By a lattice coset or simply a coset, we mean a coset $L+\nu$ in $V/L$, where $\nu\in V$. The smallest integer $a$ such that $a\nu\in L$ is called the conductor of $L+\nu$. For example, let
$L={\mathbb{Z}}e_{1}+{\mathbb{Z}}e_{2}+{\mathbb{Z}}e_{3}\cong\langle m^{2},m^{2},m^{2}\rangle\quad\text{and}\quad\nu=\frac{1}{m}(e_{1}+e_{2}+e_{3}),$
where $m\in{\mathbb{N}}$. Then the conductor of $L+d\nu$ is $m$ for any $d\in{\mathbb{Z}}$ with $(m,d)=1$. Moreover, note that any vector $y\in L+d\nu$ is written as $y=x+d\nu$ with $x=x_{1}e_{1}+x_{2}e_{2}+x_{3}e_{3}\in L$ and
$Q(y)=Q(x+d\nu)=Q\left(\sum_{i=1}^{3}\left(x_{i}+\frac{d}{m}\right)e_{i}\right)=\sum_{i=1}^{3}m^{2}\left(x_{i}+\frac{d}{m}\right)^{2}=\sum_{i=1}^{3}\left(mx_{i}+d\right)^{2}.$
Thus, writing $n$ as a sum of three squares in ${\mathcal{A}}_{d,m}^{2}$ is equivalent to finding a vector $y$ in $L+d\nu$ with $Q(y)=n$. We note that $L+\nu=L$ if and only if $\nu\in L$, equivalently, if the conductor of $L+\nu$ is $1$.
2.3. Algebraic structures of lattice cosets
Our approach requires an understanding of algebraic structures of lattice cosets: proper classes, proper genera, and proper spinor genera. To introduce these concepts, let us denote the orthogonal group $O(V)$ and the special orthogonal group $O^{+}(V)$ of $V$ by
$O(V):=\{\sigma\in\mathrm{GL}(V)\mid B(\sigma x,\sigma y)=B(x,y)\text{ for all }x,y\in V\}\quad\text{and}\quad O^{+}(V):=O(V)\cap\mathrm{SL}(V).$
Let $\theta:O^{+}(V)\rightarrow{\mathbb{Q}}^{\times}/({\mathbb{Q}}^{\times})^{2}$ be the spinor norm map (cf. [17, $\S 55$]) and denote its kernel by
$O^{\prime}(V)=\{\sigma\in O^{+}(V):\theta(\sigma)=1\}.$
Let $O_{A}^{+}(V)$ and $O_{A}^{\prime}(V)$ be the adélizations of $O^{+}(V)$ and $O^{\prime}(V)$, respectively. Define
$O^{+}(L+\nu)=\{\sigma\in O^{+}(V):\sigma(L+\nu)=L+\nu\}\quad\text{and}\quad o^{+}(L+\nu)=|O^{+}(L+\nu)|.$
Similarly, the groups $O^{+}(L_{p}+\nu)$ for any prime $p$ and $O_{A}^{+}(L+\nu)$ are defined analogously.
Let $X=L+\nu$ be a coset. The proper class $\text{cls}^{+}(X)$ of $X$ is defined as the orbit of $X$ under the action of $O^{+}(V)$. By [7, Lemma 4.2], the adelic group $O_{A}^{+}(V)$ acts on $X$, and the orbit of $X$ under this action $O_{A}^{+}(V)$ is called the proper genus of $X$, denoted by $\text{gen}^{+}(X)$. The proper spinor genus $\text{spn}^{+}(X)$ of $X$ is defined to be the orbit of $X$ under the action of the group $O^{+}(V)O_{A}^{\prime}(V)$. Clearly, we have
$\text{cls}^{+}(X)\subseteq\text{spn}^{+}(X)\subseteq\text{gen}^{+}(X).$
Moreover, [22, Proposition 2.5] shows that the number of proper spinor genera in $\text{gen}^{+}(L+\nu)$ is given by
$\left[I_{\mathbb{Q}}:{\mathbb{Q}}^{\times}\prod\limits_{p\in\Omega}\theta\left(O^{+}({L_{p}+\nu})\right)\right],$ (2.3)
where $I_{\mathbb{Q}}$ is the idèle group and $\Omega$ is the set of all places of $\mathbb{Q}$}, including the infinite place $\infty$.
2.4. Modular forms
We briefly introduce modular forms of half-integral weight below. For further details, we refer the reader to [18]. A function $f:\mathbb{H}{\,\rightarrow\,}\mathbb{C}$ satisfies a modularity of weight $\kappa\in\frac{1}{2}+{\mathbb{Z}}$ on a congruence subgroup $\Gamma\subseteq\Gamma_{0}(4)$ with character $\chi$ if $f|_{\kappa}=\chi(d)f$ for every $\gamma=\left(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right)\in\Gamma$. Here the slash operator is defined by
$f|_{\kappa}\gamma(z)=\left(\frac{c}{d}\right)\varepsilon_{d}^{2\kappa}(cz+d)^{-\kappa}f(\gamma z),$
where $\left(\frac{c}{d}\right)$ denotes the Kronecker symbol and $\varepsilon_{d}=1$ if $d\equiv 1\pmod{4}$, $\varepsilon_{d}=i$ if $d\equiv 3\pmod{4}$. We call $f(z)$ a holomorphic modular form if it satisfies the modularity condition, is holomorphic on $\mathbb{H}$, and grows at most polynomially in $y$ as $z=x+iy\rightarrow{\mathbb{Q}}\cup\{i\infty\}$. If, in addition, $f(z)\rightarrow 0$ as $z\rightarrow{\mathbb{Q}}\cup\{i\infty\}$, then we call $f$ a cusp form. If $\left(\begin{smallmatrix}1&1\\ 0&1\end{smallmatrix}\right)\in\Gamma$, then a modular form $f(z)$ on $\Gamma$ admits the following Fourier expansion:
$f(z)=\sum_{n\geq 0}a_{f}(n)q^{n},\quad\text{where }q:=e^{2\pi iz}.$
The space of modular forms (resp. cusp forms) of weight $\kappa$, character $\chi$, and congruence subgroup $\Gamma$ is denoted by $M_{\kappa}(\Gamma,\chi)$ (resp. $S_{\kappa}(\Gamma,\chi)$). The space of Eisenstein series, denoted by $E_{\kappa}(\Gamma,\chi)$, is defined as the orthogonal complement of $S_{\kappa}(\Gamma,\chi)$ in $M_{\kappa}(\Gamma,\chi)$ with respect to the Petersson inner product.
2.5. Theta series for cosets
Let $X=L+\nu$ be a coset on a quadratic space $V$ of rank $k$, and assume throughout that $B(X,X)\subseteq{\mathbb{Z}}$. For a positive integer $n$, we define
$R_{X}(n):=\{x\in X:Q(x)=n\}\quad\text{and}\quad r_{X}(n):=|R_{X}(n)|,$
and the theta series $\Theta_{X}(z)$ of the coset $X$ is defined as
$\Theta_{X}(z):=\sum\limits_{x\in X}q^{Q(x)}=\sum\limits_{n=0}^{\infty}r_{X}(n)q^{n}.$
Using [20, Proposition 2.1], one may show that $\Theta_{X}$ is a modular form of weight $\frac{k}{2}$ when $X$ is of rank $k$.
The theta series $\Theta_{\text{gen}^{+}(X)}(z)$ of $\text{gen}^{+}(X)$ and its Fourier coefficients $r_{\text{gen}^{+}(X)}(n)$ are defined by
$\Theta_{\text{gen}^{+}(X)}(z)=\sum\limits_{n=0}^{\infty}r_{\text{gen}^{+}(X)}(n)q^{n}:=\left(\sum\limits_{Y\in\text{gen}^{+}(X)}\frac{1}{o^{+}(Y)}\right)^{-1}\cdot\left(\sum\limits_{Y\in\text{gen}^{+}(X)}\frac{\Theta_{Y}(z)}{o^{+}(Y)}\right).$ (2.4)
In the same manner, we define the theta series $\Theta_{\text{spn}^{+}(X)}(z)$ of $\text{spn}^{+}(X)$ and $r_{\text{spn}^{+}(X)}(n)$ by
$\Theta_{\text{spn}^{+}(X)}(z)=\sum\limits_{n=0}^{\infty}r_{\text{spn}^{+}(X)}(n)q^{n}:=\left(\sum\limits_{Y\in\text{spn}^{+}(X)}\frac{1}{o^{+}(Y)}\right)^{-1}\cdot\left(\sum\limits_{Y\in\text{spn}^{+}(X)}\frac{\Theta_{Y}(z)}{o^{+}(Y)}\right).$ (2.5)
Here, the summations run over a complete set of representatives of proper classes in $\text{gen}^{+}(X)$ or $\text{spn}^{+}(X)$, respectively.
3. Almost square universality of ${\mathcal{A}}_{d,m}$
In this section, we discuss the almost square universality of ${\mathcal{A}}_{d,m}$ and prove Theorem 1.1. We first investigate necessary conditions and establish effective almost square universality.
Proposition 3.1.
Let $m,d\in{\mathbb{N}}$ with $d\leq m$ and $(m,d)=1$. For each $n\in{\mathbb{N}}$, write $n=mt+r_{0}d^{2}$ with unique integers $t$ and $0\leq r_{0}<m$. Define $M=1$ if $m\equiv 1\ (\mathrm{mod}\ 2)$; $M=4$ if $m\equiv 2\ (\mathrm{mod}\ 4)$; and $M=2$ if $m\equiv 0\ (\mathrm{mod}\ 4)$.
  1. (1)
    If $n$ can be written as a sum of $r$ squares in ${\mathcal{A}}_{d,m}^{2}$, then $t\equiv\frac{r-r_{0}}{m}\ (\mathrm{mod}\ M)$.
  2. (2)
    Every $n\geq N$ can be expressed as a sum of at most $Mm+3$ squares for ${\mathcal{A}}_{d,m}$, where
    $N=\begin{cases}\frac{25}{4}m^{4}+Md^{2}&\text{if }m\equiv 1\ (\mathrm{mod}\ 2),\\ \frac{m^{4}}{16}+Md^{2}&\text{if }m\equiv 2\ (\mathrm{mod}\ 4),\\ \frac{m^{4}}{4}+Md^{2}&\text{if }m\equiv 0\ (\mathrm{mod}\ 4).\end{cases}$
  3. (3)
    If ${\mathcal{A}}_{d,m}$ is $s$-almost square universal, then
    $s\geq\begin{cases}m+3&\text{if }m\equiv 1\ (\mathrm{mod}\ 2),\\ Mm+2&\text{if }m\equiv 0\ (\mathrm{mod}\ 2).\end{cases}$
Proof.
(1) We first investigate necessary conditions for $n$ to be expressed as a sum of squares in ${\mathcal{A}}_{d,m}^{2}$. Suppose that $n$ is written as below and note that
$n=\sum_{i=1}^{r}(mx_{i}+d)^{2}\quad\iff\quad mt+r_{0}d^{2}=m\left(m\sum_{i=1}^{r}x_{i}^{2}+2d\sum_{i=1}^{r}x_{i}\right)+rd^{2}.$ (3.1)
Therefore, we necessarily have $r\equiv r_{0}\ (\mathrm{mod}\ m)$, and hence (3.1) is equivalent to
$t-\frac{r-r_{0}}{m}\cdot d^{2}=m\sum_{i=1}^{r}x_{i}^{2}+2d\sum_{i=1}^{r}x_{i}.$ (3.2)
Writing $\sum_{i=1}^{r}x_{i}^{2}=a$ and $\sum_{i=1}^{r}x_{i}=b$, we have
$a\equiv b\ (\mathrm{mod}\ 2)\quad\text{and}\quad t-\frac{r-r_{0}}{m}\cdot d^{2}=ma+2db.$
Noting that $(m,d)=1$ and that $d\equiv 1\ (\mathrm{mod}\ 2)$ when $m\equiv 0\ (\mathrm{mod}\ 2)$, it follows that
$\{ma+2db\mid a,b\in{\mathbb{Z}},\ a\equiv b\ (\mathrm{mod}\ 2)\}=M{\mathbb{Z}}.$
Therefore, we necessarily have $t\equiv\frac{r-r_{0}}{m}\cdot d^{2}\equiv\frac{r-r_{0}}{m}\ (\mathrm{mod}\ M)$.
(2) For $n\geq N$, we will show that there exist integers $a,b\in{\mathbb{Z}}$ and $4\leq r\leq Mm+3$ satisfying
[…]
Then, since $a$ and $b$ satisfy (2.2), by Lemma 2.1 there exist $x_{1},x_{2},x_{3},x_{4}\in{\mathbb{Z}}$ such that $\sum_{i=1}^{4}x_{i}^{2}=a$ and $\sum_{i=1}^{4}x_{i}=b$. Hence setting $x_{5}=\cdots=x_{r}=0$, equations (3.1) and (3.2) show that $n=\sum_{i=1}^{r}(mx_{i}+d)^{2}$.
We choose $r\in{\mathbb{N}}$ to be the unique $4\leq r\leq Mm+3$ satisfying $t\equiv\frac{r-r_{0}}{m}\ (\mathrm{mod}\ M)$, which is necessary by part (1) of the proposition. To explain the choice of $a$ and $b$, for simplicity, let us set
$\tilde{t}:=t-\frac{r-r_{0}}{m}d^{2}.$
(Case 1) $m\equiv 1\ (\mathrm{mod}\ 2)$. Note that if $(a,b)=(a_{0},b_{0})$ is an arbitrary integer solution to $\tilde{t}=ma+2db$, then all integer solutions $(a,b)$ to this equation are given by
$(a_{k},b_{k})=(a_{0}-2dk,b_{0}+mk)\text{ with }k\in{\mathbb{Z}}.$
We claim that $(a,b)=(a_{k},b_{k})$ satisfies the second and third conditions in (3.3) for some $0\leq k\leq 9$. If this claim holds, then we may find $(a,b)$ satisfying (3.3) as follows: since $a=\frac{\tilde{t}-2db}{m}$, we have
$4a-b^{2}=4\cdot\frac{\tilde{t}-2db}{m}-b^{2}\geq 0$ $\iff mb^{2}+8db-4\tilde{t}\leq 0$
$\iff b\in\left[-\frac{4d}{m}-\frac{2\sqrt{4d^{2}+m\tilde{t}}}{m},-\frac{4d}{m}+\frac{2\sqrt{4d^{2}+m\tilde{t}}}{m}\right].$ (3.4)
Since $n=mt+r_{0}d^{2}=m\tilde{t}+rd^{2}\geq\frac{25}{4}m^{4}+Md^{2}$, we have
$\tilde{t}\geq\frac{25}{4}m^{4}+\frac{Mm-r}{m}d^{2}\geq\frac{25}{4}m^{4}-\frac{4d^{2}}{m},$
which implies that the length of the interval in (3.4) is at least $10m$. Therefore, there are at least ten consecutive solutions $(a,b)$ to $\tilde{t}=ma+2db$ with $4a-b^{2}\geq 0$. Among these, one can find a pair $(a,b)=(a_{k},b_{k})$ that satisfies (3.3), as claimed.
Now we prove the claim. Let $i\in\{0,1\}$ be such that $a_{i}\equiv b_{i}\ (\mathrm{mod}\ 2)$, and note that for all $j\in{\mathbb{Z}}$, we have $a_{i+2j}\equiv b_{i+2j}\ (\mathrm{mod}\ 2)$. Define $D_{k}=4a_{k}-b_{k}^{2}$. If $a_{i}\equiv b_{i}\equiv 1\ (\mathrm{mod}\ 2)$, then $D_{i}\equiv 3\ (\mathrm{mod}\ 8)$, so we may take $k=i$. Assume instead that $a_{i}\equiv b_{i}\equiv 0\ (\mathrm{mod}\ 2)$ and write $a_{i}=2a_{i}^{\prime}$ and $b_{i}=2b_{i}^{\prime}$. Then $D_{i}\equiv 0\ (\mathrm{mod}\ 4)$, and we have
[…]
In these cases, we may also take $k=i$. The remaining cases are when $b_{i}^{\prime}\equiv 1\ (\mathrm{mod}\ 2)$ with $a_{i}^{\prime}\equiv 0\ (\mathrm{mod}\ 4)$, and when $b_{i}^{\prime}\equiv 0\ (\mathrm{mod}\ 2)$ with $a_{i}^{\prime}\equiv 0\ (\mathrm{mod}\ 2)$. In the former case, note that $D_{i}/4\equiv 7\ (\mathrm{mod}\ 8)$. Since
$\frac{D_{i+2j}}{4}=2a_{i+2j}^{\prime}-b_{i+2j}^{\prime 2}=\frac{D_{i}}{4}-\left[(4d+2b_{i}^{\prime}m)j+m^{2}j^{2}\right]$
and $4d+2b_{i}^{\prime}m\equiv 2\ (\mathrm{mod}\ 4)$, we have $D_{i+2j_{0}}/4\equiv 3\ (\mathrm{mod}\ 8)$ for $j_{0}\in\{1,3\}$ with $4d+2b_{i}^{\prime}m\equiv 2j_{0}\ (\mathrm{mod}\ 8)$. Thus, we may take $k=i+2j_{0}$. For the latter case, since $b_{i+2}^{\prime}\equiv 1\ (\mathrm{mod}\ 2)$, repeating the above argument by replacing $i$ with $i+2$ shows that $(a_{k},b_{k})$ for some $k\in\{i+2,i+4,i+8\}$ satisfies (3.3). This completes the proof of the claim.
(Case 2) $m\equiv 0\ (\mathrm{mod}\ 2)$. The strategy is the same as in the previous case with appropriate modifications. Note that $\tilde{t}\equiv 0\ (\mathrm{mod}\ M)$ and $d\equiv 1\ (\mathrm{mod}\ 2)$. Suppose $(a,b)=(a_{0},b_{0})$ is an arbitrary integer solution to $\frac{\tilde{t}}{2}=\frac{m}{2}a+db$. Then all integer solutions $(a,b)$ are given by
$(a_{k},b_{k})=(a_{0}-dk,b_{0}+\frac{m}{2}k)\text{ with }k\in{\mathbb{Z}}.$
We claim that there exists some integer $0\leq k\leq M-1$ such that $(a,b)=(a_{k},b_{k})$ satisfies the second and third conditions in (3.3).
If $m\equiv 2\ (\mathrm{mod}\ 4)$, then $\frac{\tilde{t}}{2}\equiv 0\ (\mathrm{mod}\ 2)$. Hence for all $j$, we have $a_{j}\equiv b_{j}\ (\mathrm{mod}\ 2)$, and there exists some $i\in\{0,1\}$ such that $a_{i}\equiv b_{i}\equiv 1\ (\mathrm{mod}\ 2)$. For this $i$, it follows that $4a_{i}-b_{i}^{2}\equiv 3\ (\mathrm{mod}\ 8)$, so we may take $k=i$.
If $m\equiv 0\ (\mathrm{mod}\ 4)$, then let $i\in\{0,1\}$ be such that $a_{i+2j}\equiv b_{i+2j}\ (\mathrm{mod}\ 2)$ for all integers $j$. If either $a_{i}\equiv b_{i}\equiv 1\ (\mathrm{mod}\ 2)$, or $a_{i}=2a_{i}^{\prime}$ and $b_{i}=2b_{i}^{\prime}$ satisfy one of the cases in (3.5), then we may take $k=i$. Otherwise, we consider two remaining cases: either $b_{i}^{\prime}\equiv 1\ (\mathrm{mod}\ 2)$ with $a_{i}^{\prime}\equiv 0\ (\mathrm{mod}\ 4)$; or $b_{i}^{\prime}\equiv 0\ (\mathrm{mod}\ 2)$ with $a_{i}^{\prime}\equiv 0\ (\mathrm{mod}\ 2)$. In the first case, since $D_{i}/4\equiv 7\ (\mathrm{mod}\ 8)$, we have
$D_{i+2}/4=D_{i}/4-\left[(2d+b_{i}^{\prime}m)+\left(\frac{m}{2}\right)^{2}\right]\equiv 3-2\equiv 1\ (\mathrm{mod}\ 4).$
In the second case, we have $a_{i+2}^{\prime}=a_{i}^{\prime}-d\equiv 1\ (\mathrm{mod}\ 2)$ and $b_{i+2}^{\prime}=b_{i}^{\prime}+\frac{m}{2}\equiv 0\ (\mathrm{mod}\ 2)$. Therefore, we have $D_{i+2}/4=2a_{i+2}^{\prime}-b_{i+2^{\prime}}^{2}\equiv 2\ (\mathrm{mod}\ 4)$. Thus, in either case, we may take $k=i+2$. This completes the proof of the claim.
Note that from $n=m\tilde{t}+rd^{2}\geq\frac{m^{4}}{M^{2}}+Md^{2}$, we obtain
$\tilde{t}\geq\frac{m^{4}}{M^{2}}+\frac{Mm-r}{m}d^{2}>\frac{m^{4}}{M^{2}}-\frac{4d^{2}}{m}.$
Hence the length of the interval in (3.4) is greater than $M\cdot\frac{m}{2}$. Therefore, there are at least $M$ consecutive solutions $(a,b)$ to $\frac{\tilde{t}}{2}=\frac{m}{2}a+db$ with $4a-b^{2}>0$, among which one can find $(a,b)=(a_{k},b_{k})$ satisfying (3.3) by the claim. This completes the proof of part (2).
(3) First, suppose $m\equiv 0\ (\mathrm{mod}\ 2)$. Let $p$ be a prime such that $p\equiv 3\ (\mathrm{mod}\ 4)$ and $(p,mM)=1$. Consider a positive integer $n=mMt_{0}+2d^{2}$, where $t_{0}$ is a positive integer such that
$mMt_{0}\equiv-2d^{2}\ (\mathrm{mod}\ p)\quad\text{and}\quad mMt_{0}\not\equiv-2d^{2}\ (\mathrm{mod}\ p^{2}).$
Since $\text{ord}_{p}(n)=1$, the integer $n$ is not representable as a sum of two squares of integers. On the other hand, by part (1), to write $n=mMt_{0}+2d^{2}$ as a sum of $r$ squares in ${\mathcal{A}}_{d,m}^{2}$, it is necessary that $r\equiv 2\ (\mathrm{mod}\ Mm)$. Thus $r\geq mM+2$. Since there are infinitely many such $n$, we conclude that $\mathrm{ASU}(\mathcal{A}_{d,m})\geq Mm+2$.
Now suppose $m\equiv 1\ (\mathrm{mod}\ 2)$. Consider a positive integer $n=mt_{0}+3d^{2}$ such that $n\equiv 7\ (\mathrm{mod}\ 8)$. Then $n$ is not representable as a sum of three squares of integers. By part (1), in order to write this $n$ as a sum of $r$ squares in ${\mathcal{A}}_{d,m}^{2}$, it is necessary that $r\geq m+3$. Since there are infinitely many such $n$, we conclude that $\mathrm{ASU}(\mathcal{A}_{d,m})\geq m+3$. ∎
As a direct consequence of Proposition 3.1, we obtain the following corollary concerning $\mathrm{ASU}(\mathcal{A}_{d,m})$.
Corollary 3.2.
Let $m\in{\mathbb{N}}$ and $d\in{\mathbb{Z}}$ be such that $(m,d)=1$. Then
$\mathrm{ASU}(\mathcal{A}_{d,m})=\begin{cases}m+3&\text{if }m\equiv 1\ (\mathrm{mod}\ 2),\\ Mm+2\text{ or }Mm+3&\text{if }m\equiv 0\ (\mathrm{mod}\ 2).\end{cases}$
Moreover, for $m\equiv 0\ (\mathrm{mod}\ 2)$, we have $\mathrm{ASU}(\mathcal{A}_{d,m})=Mm+2$ if and only if every sufficiently large integer $n\equiv 3d^{2}\ (\mathrm{mod}\ m)$ can be written as a sum of three squares in ${\mathcal{A}}_{d,m}^{2}$.
According to Corollary 3.2, determining $\mathrm{ASU}(\mathcal{A}_{d,m})$ reduces to the case when $m$ is even and to expressing every sufficiently large integer $n\equiv 3d^{2}\ (\mathrm{mod}\ m)$ as a sum of three squares in ${\mathcal{A}}_{d,m}$. We address this problem by studying representations of integers by ternary lattice cosets.
Theorem 3.3.
Let $m\in 2{\mathbb{N}}$ and $d\in{\mathbb{Z}}$ with $(m,d)=1$. Let $L={\mathbb{Z}}e_{1}+{\mathbb{Z}}e_{2}+{\mathbb{Z}}e_{3}\cong\langle m^{2},m^{2},m^{2}\rangle$ be a ternary $\mathbb{Z}$}-lattice and let $\nu=\frac{1}{m}(e_{1}+e_{2}+e_{3})\in{\mathbb{Q}}L$. Consider the lattice coset $X^{d}=L+d\nu$.
  1. (1)
    If $m\equiv 2\ (\mathrm{mod}\ 4)$ or $m\equiv 0\ (\mathrm{mod}\ 4)$ with $3\nmid m$, then $\text{spn}^{+}(X^{d})=\text{gen}^{+}(X^{d})$.
  2. (2)
    If $m\equiv 0\ (\mathrm{mod}\ 12)$, then there are two proper spinor genera in $\text{gen}^{+}(X^{d})$.
Proof.
Recall from (2.3) that the number of proper spinor genera in $\text{gen}^{+}(X^{d})$ is given by
$\left[I_{\mathbb{Q}}:{\mathbb{Q}}^{\times}\prod\limits_{p\in\Omega}\theta\left(O^{+}({X_{p}^{d}})\right)\right].$ (3.6)
Hence, it remains to compute $\theta\left(O^{+}({X_{p}^{d}})\right)$ for all primes $p$. In fact, these can be obtained from the previously calculated local spinor norm groups for lattice cosets considered in [1] and [11]. Therefore, instead of providing an explicit computation of $\theta\left(O^{+}({X_{p}^{d}})\right)$, we explain how they can be deduced from those results.
First note that since $(m,d)=1$, [13, Lemma 3.1 (1)] implies that $O^{+}(L_{p}+d\nu)=O^{+}(L_{p}+\nu)$ for all primes $p$. Therefore, for each prime $p$, the spinor norm group $\theta\left(O^{+}({X_{p}^{d}})\right)$ is independent of $d$ as long as $(m,d)=1$.
Assume $m\equiv 2\ (\mathrm{mod}\ 4)$ and $p=2$. In this case, we have $L_{2}={\mathbb{Z}}_{2}e_{1}+{\mathbb{Z}}_{2}e_{2}+{\mathbb{Z}}_{2}e_{3}\cong\langle 4,4,4\rangle$ and
$X_{2}=L_{2}+\nu=\frac{1}{m}(e_{1}+e_{2}+e_{3})=L_{2}+\frac{1}{2}(e_{1}+e_{2}+e_{3}).$
By [14, Proposition 2.3], it follows that $O^{+}(X_{2})=O^{+}(L_{2})$, and hence $\theta\left(O^{+}({X_{2}^{d}})\right)={\mathbb{Z}}_{2}({\mathbb{Q}}_{2}^{\times})^{2}$ as described in the proof of [1, Proposition 4.2].
Otherwise, either $p$ is an odd prime or $p=2$ with $m\equiv 0\ (\mathrm{mod}\ 4)$. In this case, the spinor norm group $\theta\left(O^{+}({X_{p}^{d}})\right)$ has already been computed in [11, Proposition 3.4] under different setting. To explain this, consider the lattice coset $K+\mu$ defined by
$K=K_{(m^{\prime})}={\mathbb{Z}}f_{1}+{\mathbb{Z}}f_{2}+{\mathbb{Z}}f_{3}\cong\langle(m^{\prime}-2)^{2},(m^{\prime}-2)^{2},(m^{\prime}-2)^{2}\rangle\quad\text{and}\quad\mu:=\frac{m^{\prime}-4}{2(m^{\prime}-2)}(f_{1}+f_{2}+f_{3})$
as in [11, (2.1) and (2.2)]. By setting $m^{\prime}=m+2$, we observe that $K+\mu=L+\left(\frac{m-2}{2}\right)\nu$, and $(m,\frac{m-2}{2})=1$ since $m\equiv 0\ (\mathrm{mod}\ 4)$. Therefore, for all $d$ with $(m,d)=1$, we have $\theta\left(O^{+}({X_{p}^{d}})\right)=\theta\left(O^{+}({K+\mu})\right)$. Note also that
[…]
so the conditions for $m^{\prime}$ in [11, Proposition 3.4] correspond exactly to the condition on $m$ here. Hence,
$\theta(O^{+}(X^{d}_{p}))=\begin{cases}({\mathbb{Q}}_{3}^{\times})^{2}&\text{if }p=3\text{ and }p\mid m,\\ {\mathbb{Z}}_{p}^{\times}({\mathbb{Q}}_{p}^{\times})^{2}&\text{otherwise}.\end{cases}$
Furthermore, when $m\equiv 0\ (\mathrm{mod}\ 4)$, the index in (3.6), which counts the number of proper spinor genera in $\text{gen}^{+}(X^{d})$, can also be computed as in the proof of [11, Proposition 3.4]. This completes the proof for the case when $m\equiv 0\ (\mathrm{mod}\ 4)$.
When $m\equiv 2\ (\mathrm{mod}\ 4)$, the above computations give
$\theta\left(O^{+}({X_{p}^{d}})\right)=\begin{cases}{\mathbb{Z}}_{2}({\mathbb{Q}}_{2}^{\times})^{2}&\text{if }p=2,\\ ({\mathbb{Q}}_{3}^{\times})^{2}&\text{if }p=3\text{ and }3\mid m,\\ {\mathbb{Z}}_{p}^{\times}({\mathbb{Q}}_{p}^{\times})^{2}&\text{otherwise}.\end{cases}$
The index (3.6) equals 11 as computed in the proof of [1, Theorem 4.3], completing the proof. ∎
Using the decomposition of theta series into three parts, we obtain the following theorem.
Theorem 3.4.
Let $m\in{\mathbb{N}}$ be an even integer such that $12\nmid m$, and let $d\in{\mathbb{Z}}$ satisfy $1\leq d\leq m$ with $(m,d)=1$. Then every sufficiently large integer $n$ with $n\equiv 3d^{2}\ (\mathrm{mod}\ m)$ can be represented in the form
$n=\sum_{i=1}^{3}(mx_{i}+d)^{2}$ (3.7)
for some integers $x_{i}\in{\mathbb{Z}}$.
Proof.
Let $L={\mathbb{Z}}e_{1}+{\mathbb{Z}}e_{2}+{\mathbb{Z}}e_{3}\cong\langle m^{2},m^{2},m^{2}\rangle$ and $\nu=\frac{1}{m}(e_{1}+e_{2}+e_{3})\in{\mathbb{Q}}L$. Consider $X^{d}=L+d\nu$. Recall that $n$ can be written as in (3.7) if and only if $n$ is represented by $X^{d}$. Since $\text{gen}^{+}(X^{d})=\text{spn}^{+}(X^{d})$ by Theorem 3.3, and recalling definitions (2.4) and (2.5), we have the decomposition
$\Theta_{X^{d}}=\Theta_{\text{gen}^{+}(X^{d})}+\Theta_{X^{d}}-\Theta_{\text{gen}^{+}(X^{d})}=\Theta_{\text{gen}^{+}(X^{d})}+\Theta_{X^{d}}-\Theta_{\text{spn}^{+}(X^{d})}.$
By [13, Theorem 6.2], the difference $\Theta_{X^{d}}-\Theta_{\text{spn}^{+}(X^{d})}$ lies in the subspace of cusp forms orthogonal to unary theta functions. Hence, by Duke’s bound [9], the absolute value of the Fourier coefficients of $\Theta_{X^{d}}-\Theta_{\text{spn}^{+}(X^{d})}$ satisfy the estimate $\ll_{m,\varepsilon}n^{3/7+\varepsilon}$. On the other hand, the Fourier coefficients $r_{\text{gen}^{+}(X^{d})}(n)$ of $\Theta_{\text{gen}^{+}(X^{d})}$ satisfy the lower bound $r_{\text{gen}^{+}(X^{d})}(n)\gg_{m,\varepsilon}n^{\frac{1}{2}-\varepsilon}$, since $n$ is locally represented by $X^{d}$ and the order $\text{ord}_{p}(n)=0$ for every prime divisor $p\mid m$. Comparing these asymptotic behaviors, we conclude the proof of theorem. ∎
Theorem 3.5 ([11, Theorem 5.2]).
$\mathrm{(1)}$ If $\ell\equiv 1\ (\mathrm{mod}\ 12)$ is an odd prime, then
$x^{2}+y^{2}+z^{2}=3\ell^{2}$
has no solution in $x,y,z\in{\mathbb{Z}}$ with $x\equiv y\equiv z\equiv 5\ (\mathrm{mod}\ 12)$.
$\mathrm{(2)}$ If $\ell\equiv 7\ (\mathrm{mod}\ 12)$ is an odd prime, then
$x^{2}+y^{2}+z^{2}=3\ell^{2}$
has no solution in $x,y,z\in{\mathbb{Z}}$ with $x\equiv y\equiv z\equiv 1\ (\mathrm{mod}\ 12)$.
With all the necessary results established, we are ready to complete the proof of Theorem 1.1.
proof of Theorem 1.1.
By Corollary 3.2 and Theorem 3.4, we have
[…]
Thus, it remains to show that $\mathrm{ASU}(\mathcal{A}_{d,m})=2m+3$ if $m\equiv 0\ (\mathrm{mod}\ 12)$.
Write $m=12m_{0}$ with $m_{0}\in{\mathbb{N}}$, and let $d_{0}\in\{1,5\}$ be such that $d\equiv\pm d_{0}\ (\mathrm{mod}\ 12)$. Consider an odd prime $\ell$ satisfying
[…]
Assume that $n=mt+3d^{2}=3\ell^{2}$ and that $n$ is a sum of three squares in ${\mathcal{A}}_{d,m}^{2}$. Then
$3\ell^{2}=n=\sum_{i=1}^{3}(mx_{i}+d)^{2}=(12m_{0}x_{1}+d)^{2}+(12m_{0}x_{2}+d)^{2}+(12m_{0}x_{3}+d)^{2}.$ (3.9)
for some $x_{i}\in{\mathbb{Z}}$. By changing the sign of all $x_{i}$ if necessary, the equation $x^{2}+y^{2}+z^{2}=3\ell^{2}$ has a solution with $x\equiv y\equiv z\equiv d_{0}\ (\mathrm{mod}\ 12)$, which contradicts Theorem 3.5.
Hence, by Corollary 3.2, it suffices to show that there are infinitely many odd primes $\ell$ satisfying (3.8) and $3\ell^{2}\equiv 3d^{2}\ (\mathrm{mod}\ m)$, equivalently, $\ell^{2}\equiv d^{2}\ (\mathrm{mod}\ 4m_{0})$. Write $m_{0}=2^{a}3^{b}m_{1}$ with $(m_{1},6)=1$. Then by the Chinese Remainder Theorem and Hensel’s Lemma, there exists $\ell$ such that $\ell\equiv d\ (\mathrm{mod}\ m_{1})$ and
[…]
Such $\ell$ satisfies $\ell^{2}\equiv d^{2}\ (\mathrm{mod}\ 4m_{0})$. By Dirichlet’s theorem on arithmetic progression. This concludes the proof that $\mathrm{ASU}({\mathcal{A}}_{d,m})=2m+3$ when $m\equiv 0\ (\mathrm{mod}\ 12)$. ∎
4. Square universality of ${\mathcal{A}}_{d,m}$
In this section, we discuss the square universality of ${\mathcal{A}}_{d,m}$ and prove Theorem 1.2. We first investigate the necessary conditions.
Lemma 4.1.
Let $m\in{\mathbb{N}}$ and $d\in{\mathbb{Z}}$ with $(m,d)=1$. If ${\mathcal{A}}_{d,m}$ is square universal, then we have
$d\equiv\pm 1\ (\mathrm{mod}\ m)\quad\text{and}\quad\mathrm{SU}({\mathcal{A}}_{d,m})\geq\max(m^{2}-2m,\mathrm{ASU}(\mathcal{A}_{d,m}))$
for these integers $d$.
Proof.
To represent $1$ as a sum of squares in ${\mathcal{A}}_{d,m}^{2}$, we must have either $1\in{\mathcal{A}}_{d,m}$ or $1\in{\mathcal{A}}_{d,m}$. Hence, $d\equiv\pm 1\ (\mathrm{mod}\ m)$. Moreover, to represent $(m-1)^{2}-1=m^{2}-2m$, we need at least $m^{2}-2m$ squares in ${\mathcal{A}}_{d,m}^{2}$, since $(m-1)^{2}$ is the second smallest square in ${\mathcal{A}}_{d,m}^{2}$. Therefore, $\mathrm{SU}({\mathcal{A}}_{d,m})\geq m^{2}-2m$. Since by definition $\mathrm{SU}(\mathcal{A}_{d,m})\geq\mathrm{ASU}(\mathcal{A}_{d,m})$, the claim follows. ∎
Lemma 4.2.
Let $m\geq 8$ be an integer, and let $d$ be an integer with $d\equiv\pm 1\ (\mathrm{mod}\ m)$. Then every positive integer $n\leq 2m(m-1)^{2}$ can be written as a sum of at most $m^{2}-2m$ squares in ${\mathcal{A}}_{d,m}^{2}$.
Proof.
Without loss of generality, we may assume $d=1$. Recall that the three smallest squares in ${\mathcal{A}}_{1,m}^{2}$ are $1$, $(m-1)^{2}$, and $(m+1)^{2}$. Let $k,r$ be integers such that $0\leq k\leq 2m-1$ and $k\leq r\leq m^{2}-2m$. Setting $x_{1}=\cdots=x_{k}=-1$ and $x_{k+1}=\cdots=x_{r}=0$, we have
$\sum_{i=1}^{r}(mx_{i}+1)^{2}=k\cdot(m-1)^{2}+(r-k).$
Hence, every integer $n$ such that
$k(m-1)^{2}\leq n\leq k(m-1)^{2}+(m^{2}-2m-k)=(k+1)(m-1)^{2}-(k+1)$
can be represented as a sum of squares in ${\mathcal{A}}_{1,m}^{2}$. Therefore, all integers $n\leq 2m(m-1)^{2}$ can be represented in this manner, except those $n$ satisfying
$(k+1)(m-1)^{2}-k\leq n\leq(k+1)(m-1)^{2}-1\quad\text{with }1\leq k\leq 2m-1.$ (4.1)
Now, taking $x_{1}=1$, $x_{2}=\cdots=x_{k}=-1$, and $x_{k+1}=\cdots=x_{r}=0$, we obtain
$\sum_{i=1}^{r}(mx_{i}+1)^{2}=(m+1)^{2}+(k-1)\cdot(m-1)^{2}+(r-k)=(k+1)^{2}(m-1)^{2}-(m^{2}-6m+1).$
Hence every integer $n$ in the range
$(k+1)^{2}(m-1)^{2}-(m^{2}-6m+1)\leq n\leq(k+1)^{2}(m-1)^{2}+(4m-k-1)$
is representable as a sum of squares in ${\mathcal{A}}_{1,m}^{2}$. Since $k\leq 2m-1\leq m^{2}-6m+1$ and $4m-k-1\geq 4m-(2m-1)-1\geq-1$, this range covers all integers $n$ in (4.1). This completes the proof. ∎
Although Proposition 3.1 (2) provides an effective bound for almost universality, it remains somewhat too large to guarantee that every small integer is representable as a sum of squares in ${\mathcal{A}}_{d,m}^{2}$. In the following lemma, we derive a sharper bound, from which Lemma 4.2 suffices to complete the verification of $\mathrm{SU}(\mathcal{A}_{d,m})$ for $d\equiv\pm 1\ (\mathrm{mod}\ m)$.
Lemma 4.3.
Let $m\geq 6$ be an integer, and define $M=1$ if $m\equiv 1\ (\mathrm{mod}\ 2)$, $M=4$ if $m\equiv 2\ (\mathrm{mod}\ 4)$, and $M=2$ if $m\equiv 0\ (\mathrm{mod}\ 4)$. Then every integer $n\geq 2m(m-1)^{2}$ can be written as a sum of at most $Mm+3$ squares in ${\mathcal{A}}_{d,m}^{2}$.
Proof.
Write $n=mt+r_{0}$ with unique integers $t,r_{0}\in{\mathbb{Z}}$ such that $0\leq r_{0}<m$. Let $r$ be the unique integer satisfying $m+4\leq r\leq Mm+3$ and $t\equiv\frac{r-r_{0}}{m}\ (\mathrm{mod}\ M)$. Note that the equation
$n=\sum_{i=1}^{r}(mx_{i}+1)^{2}$
has a solution $x_{i}\in{\mathbb{Z}}$ if and only if
$\tilde{t}:=t-\frac{r-r_{0}}{m}=m\left(\sum_{i=1}^{4}x_{i}^{2}\right)+2\left(\sum_{i=1}^{4}x_{i}\right)+\sum_{i=5}^{r}(mx_{i}^{2}+2x_{i}).$
Define the set
$S=\begin{cases}\{k(m-2)\mid 1\leq k\leq m\}\cup\{k(m+2)\mid 0\leq k\leq m-1\}&\text{if }m\equiv 1\ (\mathrm{mod}\ 2),\\ \{k(\frac{m}{2}+1)\mid 0\leq k\leq m-1\}&\text{if }m\equiv 0\ (\mathrm{mod}\ 2).\end{cases}$
One may verify that, with $e\in\{1,2\}$ such that $e\equiv m\ (\mathrm{mod}\ 2)$,
  1. (1)
    $S\subseteq\left\{\sum_{i=5}^{k}\frac{1}{e}(mx_{i}^{2}+2x_{i})\mid x_{i}\in\{0,\pm 1\}\right\}$,
  2. (2)
    there exists $s\in S$ such that $\frac{\tilde{t}}{e}\equiv\frac{2}{e}+s\ (\mathrm{mod}\ \frac{m}{e})$ and $\frac{\tilde{t}}{e}\not\equiv\frac{2}{e}+s\ (\mathrm{mod}\ 2\frac{m}{e})$.
Choosing $s\in S$ satisfying (2), we set
$a:=\frac{\tilde{t}/e-(2/e+s)}{m/e},$
which is an odd positive integer since
$\tilde{t}-2-es=\frac{n-r}{m}-2-es\geq 2(m-1)^{2}-\frac{4m+3}{m}-2-(m-1)(m+2)>0.$
By Lemma 2.1, there exist $x_{1},x_{2},x_{3},x_{4}\in{\mathbb{Z}}$ such that
$\begin{cases}x_{1}^{2}+x_{2}^{2}+x_{3}^{2}+x_{4}^{2}=a,\\ x_{1}+x_{2}+x_{3}+x_{4}=1.\end{cases}$
Let $x_{5},\ldots,x_{r}\in\{0,\pm 1\}$ be such that $es=\sum_{i=5}^{r}(mx_{i}^{2}+2x_{i})$. Then
$\sum_{i=1}^{r}(mx_{i}+1)^{2}=m\left(m\sum_{i=1}^{r}x_{i}^{2}+2\sum_{i=1}^{r}x_{i}\right)+r=m(ma+2+es)+r=m\tilde{t}+r=n.$
This completes the proof. ∎
Proof of Theorem 1.2.
By Lemma 4.1, we may assume that $d\equiv\pm 1\ (\mathrm{mod}\ m)$. For $m=2,3,4$, we have
$\mathrm{SU}({\mathcal{A}}_{d,2})=\mathrm{SU}({\mathcal{A}}_{d,4})=\mathrm{SU}({\mathbb{Z}}-2{\mathbb{Z}})=10\quad\text{and}\quad\mathrm{SU}({\mathcal{A}}_{d,3})=\mathrm{SU}({\mathbb{Z}}-3{\mathbb{Z}})=6\$
by [15, Propositions 2.3 and 2.4].
For $m\geq 8$, note that $m^{2}-2m\geq 4m+3$. Hence by Lemmas 4.2 and 4.3, every positive integer can be written as a sum of at most $m^{2}-2m$ squares in ${\mathcal{A}}_{d,m}^{2}$. Thus, we have $\mathrm{SU}(\mathcal{A}_{d,m})\leq m^{2}-2m$. On the other hand, By Lemma 4.1 and Theorem 1.1, $\mathrm{SU}(\mathcal{A}_{d,m})\geq\max(m^{2}-2m,\mathrm{ASU}(\mathcal{A}_{d,m}))=m^{2}-2m$. Therefore, $\mathrm{SU}(\mathcal{A}_{d,m})=m^{2}-2m$.
If $m=5$ or $m=7$, then every integer $n\geq\frac{25}{4}m^{4}+1$ is a sum of at most $m+3$ squares in ${\mathcal{A}}_{d,m}$ by Proposition 3.1. For $m=7$, a direct check shows that every positive integer $n<\frac{25}{4}\cdot 7^{4}+1=15007.25$ is a sum of at most $m^{2}-2m=35$ squares in ${\mathcal{A}}_{d,m}^{2}$, hence $\mathrm{SU}({\mathcal{A}}_{d,7})=m^{2}-2m$. On the other hand, for $m=5$, all positive integers $n\leq\frac{25}{4}\cdot 5^{4}+1=3907.25$ are sums of at most $m^{2}-2m=15$ squares in ${\mathcal{A}}_{d,m}^{2}$, except for $n=31$. Thus, $\mathrm{SU}({\mathcal{A}}_{d,5})=16$.
Finally, for $m=6$, we show that $\mathrm{SU}(\mathcal{A}_{d,m})=26$. By Lemma 4.1 and Theorem 1.1, we have
$\mathrm{SU}({\mathcal{A}}_{d,6})\geq\max(m^{2}-2m,4m+2)=26.$
On the other hand, every positive integer $n\geq 2m(m-1)^{2}=300$ is a sum of at most 27 squares in ${\mathcal{A}}_{d,6}^{2}$ by Lemma 4.3, and a direct check confirms this also holds for all $n<300$. Therefore, to prove that $\mathrm{SU}({\mathcal{A}}_{d,6})=26$, it suffices to show that every $n=mt+3$ with $t\equiv 0\ (\mathrm{mod}\ 4)$ can be written as a sum of three squares in ${\mathcal{A}}_{d,6}^{2}$.
Let $n=24t^{\prime}+3$ with $t^{\prime}\in{\mathbb{N}}$. Since $n\equiv 3\ (\mathrm{mod}\ 8)$, it is representable as a sum of three squares of odd integers. We claim that there exists a representation $n=x^{2}+y^{2}+z^{2}$ with $x,y,z\in{\mathbb{Z}}$ such that $(xyz,6)=1$. If $n\not\equiv 0\ (\mathrm{mod}\ 9)$, then any such representation satisfies $(xyz,3)=1$. If $n\equiv 0\ (\mathrm{mod}\ 9)$, then by [15, Lemma 2.2], there exists a solution $x,y,z\in{\mathbb{Z}}$ with $(xyz,3)=1$. Since all $x,y,z$ are odd, it follows that $(xyz,6)=1$. By changing signs if necessary, we may assume that $x,y,z\in{\mathcal{A}}_{d,6}$, hence $n$ is a sum of three squares in ${\mathcal{A}}_{d,6}^{2}$. This completes the proof. ∎